OpenStudy (anonymous):

Find f'(2) = u(t) dot v(t), u(2)=<1,2,-1>, u'(2)=<3,0,4>, and v(2)=. Please show work

jimthompson5910 (jim_thompson5910):

v(t)=<t,t^(2),t^(3)> v ' (t) = <1,2t, 3t^2> ... derive each component separately v ' (2) = <1,2(2), 3(2)^2> v ' (2) = <1, 4, 12>

jimthompson5910 (jim_thompson5910):

now I think you meant to say f(t) = u(t) + v(t) and you want to find f ' (2)

jimthompson5910 (jim_thompson5910):

if so, then f(t) = u(t) + v(t) f ' (t) = u ' (t) + v ' (t) f ' (2) = u ' (2) + v ' (2) f ' (2) = <3,0,4> + <1, 4, 12> I'll let you finish up

OpenStudy (anonymous):

I'm sorry, they wanted the dot product.

jimthompson5910 (jim_thompson5910):

the dot product between which two vectors u ' (2) and v ' (2) ?

OpenStudy (anonymous):

The question is strange but it wants the dot product of u(t) and v(t), then find f'(2) of that dot product's answer. The book is getting an answer of 35.

OpenStudy (anonymous):

f'(2)=u(t) dot v(t)

OpenStudy (anonymous):

The v in the question has a 2 in the parenthesis

OpenStudy (anonymous):

The mistakes in the question has been corrected.

OpenStudy (anonymous):

I tried to find the vector equation that was used to get u(2) and u'(2) but I can't figure it out.

jimthompson5910 (jim_thompson5910):

so you're essentially solving for t? this problem is so bizarre

jimthompson5910 (jim_thompson5910):

This is what I got, but I'm not sure on the notation... u(2)=<1,2,-1>, u'(2)=<3,0,4>, v(2)=<t,t^(2),t^(3)> v'(2) = <1,2t,3t^2> f(t) = u(t) dot v(t) f ' (t) = u ' (t) dot v(t) + u(t) dot v ' (t) f ' (t) = u ' (2) dot v(2) + u(2) dot v ' (2) f ' (t) = <3,0,4> dot <t,t^(2),t^(3)> + <1,2,-1> dot <1,2t,3t^2> f ' (t) = 3t+4t^3 + 1+4t-3t^2 f ' (2) = 3(2)+4(2)^3 + 1+4(2)-3(2)^2 f ' (2) = 35